OpenStudy (studio-games):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2−x and y = 4x + 3 intersect are the solutions of the equation 2−x = 4x + 3. Part B: Make tables to find the solution to 2−x = 4x + 3. Take the integer values of x only between −3 and 3. Part C: How can you solve the equation 2−x = 4x + 3 graphically?

OpenStudy (studio-games):

@hartnn

OpenStudy (studio-games):

Can someone help me please?

Nnesha (nnesha):

\[\large\rm \color{red}{y} = 2−x ~~~~~~~~~~~ \color{Red}{y} = 4x + 3\] both equations are in slope intercept y = 2-x and y also equal to 4x+3 you can replace one of the y with other equation

OpenStudy (studio-games):

I am really bad with inequalities. Can you elaborate please?

OpenStudy (studio-games):

Step by step help preferably.

Nnesha (nnesha):

give me a second

OpenStudy (studio-games):

Thank you.

Nnesha (nnesha):